OpenStudy (anonymous):

convert the decimal to a fraction. 0.35

OpenStudy (anonymous):

35/100

OpenStudy (unklerhaukus):

\[0.35=\frac{0.35}{1}\]\[\quad=\frac{3.5}{10}=\frac{35}{100}\]\[\quad=\frac{7}{20}\]

OpenStudy (anonymous):

OH SNAP YAHOO YOU JUST GOT SERVED!